Show that π‘ˆ(βˆ’π‘“, 𝑝) = βˆ’πΏ(𝑓, 𝑝) and 𝐿(βˆ’π‘“, 𝑝) = βˆ’π‘ˆ(𝑓, 𝑝).Β 


1
Expert's answer
2021-10-11T16:03:03-0400

π‘ˆ(𝑓, 𝑝) and 𝐿(𝑓, 𝑝) are upper and lower Riemann sums for the partition p.

"𝑈(𝑓, 𝑝)=\sum M_i\Delta x_i,\ L(𝑓, 𝑝)=\sum m_i\Delta x_i"

"M_i=sup\{f(x):x_{i-1}\le x \le x_i\},\ m_i=inf\{f(x):x_{i-1}\le x \le x_i\}"


"sup\{-f(x):x_{i-1}\le x \le x_i\}=-inf\{f(x):x_{i-1}\le x \le x_i\}=-m_i"

"U(-f,p)=-\sum m_i\Delta x_i=-L(f,p)"


"inf\{-f(x):x_{i-1}\le x \le x_i\}=-sup\{f(x):x_{i-1}\le x \le x_i\}=-M_i"

"L(-f,p)=-\sum M_i\Delta x_i=-U(f,p)"